OpenStudy (anonymous):

what is the intersection point of f(x)=3x-2 and g(x)=5x+6

OpenStudy (anonymous):

3x-2=5x+6; 2x=-2-6=-8; x=-4 Your answer is x=-4 Verify ... f(x)=-12-2=-14; g(x)=5*-4+6 = -14.